# Emergence of non-ergodic multifractal quantum states in geometrical fractals

### v3

@@ Methods
The multifractal nature of an eigenstate $\Phi$ can be further characterized by its singularity spectrum $f(\alpha)$. This function describes the fractal dimension of the set of sites $i$ where the local wavefunction intensity $|\langle i|\Phi\rangle|^{2}$ scales as $N^{-\alpha}$ in the large system size limit ($N\to\infty$[71, 56, 72]. While $f(\alpha)$ can, in principle, be determined from the distribution of site-specific scaling exponents:
$\alpha_{i}=-\frac{\ln(|\langle i|\Phi\rangle|^{2})}{\ln N},$
a direct computation from this definition can be numerically unstable. We therefore employ an indirect method, which relies on the generalized inverse participation ratios, $\text{IPR}_{q}(\Phi)$ as in Eq. (2), and their established scaling behavior, $\text{IPR}_{q}(\Phi)\sim N^{-\tau_{q}(\Phi)}$, characterized by the mass exponents $\tau_{q}$.
@@
From this procedure, one gets the convex hull of the singularity spectrum and the characteristic shape of the $f(\alpha)$ curve allows for the classification of states. Extended states are distinguished by an $f(\alpha)$ spectrum that, in the ideal limit, is a single point at $(\alpha=1,f(\alpha)=1)$. Numerically, this is observed as a narrow peak around this point, where the peak value is the capacity dimension $f(1)=D_{0}=1$. Localized states, with wavefunctions confined to $\mathcal{O}(1)$ sites, exhibit generalized fractal dimensions $D_{q}\approx 0$ for $q>0$, implying $\tau_{q}\approx 0$ for $q>0$. Consequently, their $f(\alpha)$ spectrum effectively collapses to the point $\alpha_{0}=0$ with $f(\alpha_{0})=0$. For these states, the capacity dimension $D_{0}=-\tau_{0}\approx 0$. Finally, multifractal states are characterized by a broader, typically convex $f(\alpha)$, independent of the size of the system and spanning a finite range of $\alpha$ values. The maximum of this curve occurs at $\alpha_{0}=\alpha_{q=0}$ and it can deviate from 1. The peak’s value is the dimension $f(\alpha_{0})=D_{0}$, with $10<D_{0}<$ 1; this is conceptually equivalent to the fractal dimension of the support set as obtained by a box-counting methodology. The overall width of the $f(\alpha)$ spectrum quantifies the degree of multifractality. The universality of certain symmetry relations for $f(\alpha)$ [72], such as $f(1+\alpha)=f(1-\alpha)+\alpha$, is restricted to specific systems or observables [74, 75] and is not generally assumed for wavefunctions in our context.
Two-eigenstates correlator
@@
[…]
Within this formalism, for extended states, the wavefunctions are considered spread over all $N$ sites of the system, and averaged quantities may scale accordingly. In the localized case, wavefunctions have their support in a finite volume proportional to the square of the characteristic localization length, and the density correlator is generally dominated – apart from a nonsingular contribution – by a pole singularity:
@@
Such two-point density correlations have been widely applied in the study of disordered quantum systems [76, 77, 56], including those lacking translational invariance [16, 78], and also to investigate intriguing phenomena such as eigenstates mutual avoidance [53]. Since the system considered in this work—a quantum system on a geometrical fractal—is deterministic and possesses no quenched disorder, we directly compute a spectral correlation function $\mathcal{F}(\omega,E)$ without ensemble averaging. The form relevant to our analysis is:
@@
This function effectively quantifies overlaps between eigenstates separated by energy $\omega$. For numerical simulations, the $\delta$-functions were smeared using Gaussian distributions.
For ideal extended states, such as those of the Gaussian Orthogonal Ensemble, $\mathcal{F}(\omega,E)=1$ across the spectral range, indicating effectively uncorrelated eigenfunctions. In more realistic extended systems, like typical metals, a broad plateau where $\mathcal{F}(\omega,E)>1$ often appears. The plateau is often bounded by a cutoff energy [79] – the Thouless energy $E_{Th}$ – at the scale of the lattice spacing, beyond which the system returns to the uncorrelated regime. As the system transitions towards localization, these correlations become more pronounced for small energy separations, causing the correlation plateau to grow in height while narrowing. This high, narrow feature is then followed by the signature exponential decay of correlations that recovers the uncorrelated regime at larger values of $\omega$.
For a multifractal state, the correlation plateau is present only within a narrow energy interval, and the Thouless energy is expected to shrink to zero in the thermodynamic limit. In this case, the uncorrelated limit at large values of $\omega$ is preceded by a power-law decay $\mathcal{F}(\omega,E)\sim\omega^{-\mu}$, where the exponent $\mu$ is related to the correlation dimension $D_{2}$ by $\mu\approx 1-D_{2}(E)$ [76, 53]. The system size $N$ influences $\mathcal{F}(\omega,E)$ for multifractal states, providing avenues to determine fractal dimensions. For instance, the height of the plateau at low energies is theoretically expected to scale as $\mathcal{F}(0,E)\propto N^{1-D_{2}(E)}$, reflecting the underlying dimension $D_{2}(E)$. For NEMs, this implies that $\mathcal{F}(\omega,E)$ for different system sizes $N$ should collapse onto a single curve when the vertical axis is rescaled by $N^{1-D_{2}(E)}$. Finite-size corrections – not considered in this work – can be taken into account to improve the extrapolation of $D_{2}(E)$ to the thermodynamic limit.
@@ Appendix A Sierpiński Gasket geometry
with $N_{g}(1)=9$ active sites in the $g=1$ building block (as illustrated in Fig. 1 of the MT).
@@ Appendix B On the Hierarchy of Spectral Gaps
the (shifted and scaled) energy levels $x_{g}$ of generation $g$ are related to those of generation $g-1$ by the recurrence relation [80]:
@@
with $\gamma=15/4$ for the standard SG case. This iterative process naturally defines a hierarchy: an initial energy interval containing the spectrum is recursively divided into subintervals, leading to the formation of a self-similar structure of bands and gaps. The system described by Eq. (15) can be approximated by a piece-wise linear model, which allows for the retrieval of the full distribution of all gap lengths, as detailed in [64].
Appendix C Amplitude profiles of representative Extended, Localized, and Non-Ergodic Multifractal States
The spatial distributions of the representative extended, localized, and non-ergodic multifractal (NEM) states, as analysed in the MT, are visualized through their wavefunction intensity profiles, in Fig. S10.
Figure S10: Intensity profiles of the representative ergodic ($E\simeq-3.667$), localized ($E\simeq-2.184$), and NEM ($E\simeq 1.509$) states. Red color represents high intensity, and black corresponds to negligibly small intensity. For the sake of simplicity, the defects that reshape the triangular lattice into a SG are shown as voids. In the ergodic phase, the intensity is homogeneously distributed across all sites, indicating delocalization. The localized state exhibits a strong concentration of intensity in a small region, in this case near the corner. The NEM state has a heterogeneous intensity profile with strong fluctuations that differs from either extended or localized states and repeating across all scales.
@@ Appendix D Other examples of NEMs
In this section, we extend our findings regarding the emergence of NEM states in the SG, by presenting additional data for several representative NEMs at different energies, all analyzed following the methodologies detailed in the MT. Specifically, their thermodynamic mass exponents $\tau(E)$ and the related singularity spectra $f(\alpha)$ are presented in Fig. S11. Furthermore, the spectral density-density correlation functions $\mathcal{F}(\omega,E)$ for these same states are displayed in Fig. S12, and their wavefunction intensity profiles are shown in Fig. S13.
Figure S11: Multifractal analysis of Non-Ergodic Multifractal (NEM) states. Left panel: Thermodynamic mass exponents $\tau(E)$ as a function of the moment order $q$ for three NEM states at energies $E\simeq-1.481$ (red circles), $E\simeq-1.287$ (orange squares), and $E\simeq 0.752$ (blue triangles). The dashed line represents the theoretical expectation for GOE statistics ($D_{q}=1$, implying $\tau(q)=q-1$), while the dotted line indicates behavior typical of localized states ($\tau(q)\simeq 0$ for $q>0$). Right panel: The singularity spectra $f(\alpha)$ for the same NEM states, illustrating their different strength of multifractality.
Figure S12: Frequency correlator $\mathcal{F}(\omega,E)$ for three distinct types of non-ergodic multifractal states, shown in comparison to the Gaussian Orthogonal Ensemble (GOE) benchmark (dashed line). While all displayed NEM correlators exhibit the general pattern of a narrow plateau at small energy separations $\omega$ followed by a power-law fall-off, variations in these characteristics distinguish the different types. Shaded area highlights the $\omega$-regime where the distinct power-law decays are most evident.
Figure S13: Intensity profiles of different types of Non-Ergodic Multifractal (NEM) states. The color scale indicates $|\psi|^{2}$ intensity, from negligible (black regions) to high (red regions).
Top row: Intensity distributions for the NEM states states detailed in Sec. D. Energies are (from left to right): $E\simeq-1.481$, $E\simeq-1.287$, and $E\simeq 0.752$. Bottom row: Other wavefunction structures at energies (from left to right): $E\simeq-3.554$, $E\simeq-1.018$, and $E\simeq-0.498$. In these supplementary examples, some states are made with “pieces of multifractals” and display properties of either insulators or metals [53], while others have spatial features reminiscent of bulk and edge modes previously identified in fractal geometries [34, 35].
@@ Appendix E Characterization of Weakly Ergodic and Special Multifractal Eigenstates
The SG eigenspectrum hosts states with features that go beyond the exemplary cases classified in the MT. This section presents a study of examples of weakly ergodic states [81] (specifically at energies $E\simeq 0.000$ and $E\simeq 2.000$) and also examines the state at $E\simeq 1.000$ that shows special properties. This latter state is remarkable for its robustness and potential for quantum control, as discussed in the MT, warranting a more detailed investigation of its characteristics here. For these states, we employ the same finite-size scaling methodology for mass exponents $\tau_{q}(E)$ and additionally examine their wavefunction amplitudes and statistical properties (such as skewness).
E.1 Mass Exponent Analysis
The finite-size scaling of the effective mass exponents $\tau_{q}(E,N)$ – calculated as per Eq. (6) from the MT – is performed for system sizes ranging from generation $g=5$ to $g=10$ and presented in Fig. S14. The thermodynamic values $\tau_{q}(E)$ are displayed in right panel of Fig. S15.
@@ E.1 Mass Exponent Analysis
Figure S14: Finite-size scaling extrapolation of mass exponents for representative states of energies: $E\simeq 0.000,E\simeq 1.000,E\simeq 2.000$. The thermodynamic value, $\tau_{q}(E)$, is obtained as the $y$-intercept of a linear fit (dashed lines) to the data points from SG generations $g=5$ through $g=10$.
Here, one can observe that while for small moment orders $q$ the mass exponents $\tau_{q}(E)$ for these states tend to follow the $\tau_{q}\approx q-1$ behavior characteristic of extended systems, deviations emerge at larger $q$. Although the thermodynamic fractal dimension $D_{2}$ of these states is high ($0.85\lesssim D_{2}\lesssim 0.95$), there are deviations—in particular for $E\simeq 1.000$—in $\tau_{q}(E)$ at $q>2$, indicating a departure from simple ergodicity and calling for a more detailed examination. In the right panel of Fig. S15, the corresponding singularity spectra.
Figure S15: Left panel: Mass exponents $\tau_{q}(E)$ as a function of moment order $q$for representative states at energies: $E\simeq 0.000$ (red circles), $E\simeq 1.000$ (orange squares), $E\simeq 2.000$ (blue triangles). Statistical errors are included; where not visible, these error bars are smaller than the marker size. The dashed black line indicates the behavior of ideal extended states ($\tau_{q}=q-1$), while the dotted black line represents ideal localized states ($\tau_{q}=0$).
Right panel: The singularity spectra for these states, showing a parabolic a pronounced parabolic shape.

### v4

@@ Methods
The multifractal nature of an eigenstate $\Phi$ can be further characterized by its singularity spectrum $f(\alpha)$. This function describes the fractal dimension of the set of sites $i$ where the local wavefunction intensity $|\langle i|\Phi\rangle|^{2}$ scales as $N^{-\alpha}$ in the large system size limit ($N\to\infty$[73, 58, 74]. In principle, $f(\alpha)$ can be determined from site-specific scaling exponents:
$\alpha_{i}=-\frac{\ln(|\langle i|\Phi\rangle|^{2})}{\ln N},$ (9)
but for finite systems this direct computation may not converge to a size-independent form. We therefore employ an indirect method, which relies on the generalized inverse participation ratios, $\text{IPR}_{q}(\Phi)$ as in Eq. (2), and their established scaling behavior, $\text{IPR}_{q}(\Phi)\sim N^{-\tau_{q}(\Phi)}$, characterized by the mass exponents $\tau_{q}$.
@@
This procedure generates the convex hull of the singularity spectrum, whose characteristic shape allows for the classification of states. Extended states are distinguished by an $f(\alpha)$ spectrum that, in the ideal limit, is a single point at $(\alpha=1,f(\alpha)=1)$. Numerically, this is observed as a narrow peak around this point, where the peak value is the capacity dimension $f(1)=D_{0}=1$. Localized states, with wavefunctions confined to $\mathcal{O}(1)$ sites, exhibit generalized fractal dimensions $D_{q}\approx 0$ for $q>0$, implying $\tau_{q}\approx 0$ for $q>0$. Consequently, their $f(\alpha)$ spectrum effectively collapses to the point $\alpha_{0}=0$ with $f(\alpha_{0})=0$. For these states, the capacity dimension $D_{0}=-\tau_{0}\approx 0$. Finally, multifractal states are characterized by a broader, typically convex $f(\alpha)$, independent of the size of the system and spanning a finite range of $\alpha$ values. The maximum of this curve occurs at $\alpha_{0}=\alpha_{q=0}$ and it can deviate from 1. The peak’s value is the dimension $f(\alpha_{0})=D_{0}$, with $0<D_{0}<1$; this is conceptually equivalent to the fractal dimension of the support set as obtained by a box-counting methodology. The overall width of the $f(\alpha)$ spectrum quantifies the degree of multifractality. The universality of certain symmetry relations for $f(\alpha)$ [74], such as $f(1+\alpha)=f(1-\alpha)+\alpha$, is restricted to specific systems or observables [76, 77] and is not generally assumed for wavefunctions in our context.
Two-eigenstate Correlator
@@
where angle brackets $\langle\dots\rangle$ denote an average over disorder realizations, $n(E)$ is the density of states, and $\omega$ is the energy difference between the two states.
Within this formalism, for extended states, the wavefunctions are considered spread over all $N$ sites of the system, and averaged quantities may scale accordingly. In the localized case, wavefunctions have their support in a finite volume proportional to the square of the characteristic localization length, and the density correlator is generally dominated— apart from a nonsingular contribution—by a pole singularity:
@@
Such two-point density correlations have been widely applied in the study of disordered quantum systems [78, 79, 58], including those lacking translational invariance [16, 80], and also to investigate intriguing phenomena such as mutual avoidance of eigenstates [55]. Since the system considered in this work—a quantum system on a geometric fractal—is deterministic and possesses no quenched disorder, we directly compute a spectral correlation function $\mathcal{F}(\omega,E)$ without ensemble averaging. The form relevant to our analysis is:
@@
This function effectively quantifies overlaps between eigenstates separated by energy $\omega$. For numerical simulations, each $\delta$-function is smeared by a normalized Gaussian, and $\mathcal{F}(\omega,E)$ is evaluated on a predefined logarithmically spaced $\omega$ grid. The Gaussian factors are applied separately to the numerator and the denominator before taking their ratio. To improve statistics, as in the IPR analysis, we average over a narrow energy window centered at $E$ by including all eigenstate pairs whose mean energy $(\lambda_{i}+\lambda_{j})/2$ lies inside that window and averaging their contributions. This local energy averaging is performed within a single deterministic spectrum and does not involve any averaging over disorder.
For ideal extended states, such as those of the Gaussian Orthogonal Ensemble, $\mathcal{F}(\omega,E)=1$ across the spectral range, indicating effectively uncorrelated eigenfunctions. In more realistic extended systems, like typical metals, a broad plateau where $\mathcal{F}(\omega,E)>1$ often appears. The plateau is often bounded by a cutoff energy [81]—the Thouless energy $E_{Th}$—at the scale of the lattice spacing, beyond which the system returns to the uncorrelated regime. As the system transitions towards localization, these correlations become more pronounced for small energy separations, causing the correlation plateau to grow in height while narrowing. This high, narrow feature is then followed by the signature exponential decay of correlations that recovers the uncorrelated regime at larger values of $\omega$.
For a multifractal state, the correlation plateau is present only within a narrow energy interval, and the Thouless energy is expected to shrink to zero in the thermodynamic limit. In this case, the uncorrelated limit at large values of $\omega$ is preceded by a power-law decay $\mathcal{F}(\omega,E)\sim\omega^{-\mu}$, where the exponent $\mu$ is related to the correlation dimension $D_{2}$ by $\mu\approx 1-D_{2}(E)$ [78, 55]. The system size $N$ influences $\mathcal{F}(\omega,E)$ for multifractal states, providing avenues to determine fractal dimensions. For instance, the plateau height at small $\omega$ is expected to scale as $\mathcal{F}(0,E)\propto N^{1-D_{2}(E)}$. More generally, scaling theory predicts that the $\mathcal{F}(\omega,E)$ curves for different $N$ should collapse under rescaling [55], however we do not apply this in the present work. Instead, we extract the decay exponents $\mu(N)$ directly from the raw $\mathcal{F}(\omega,E;N)$ data, and determine $D_{2}$ via the relation $\mu(N)\approx 1-D_{2}(N)$.
Finite-size corrections, not considered here, could be included to improve the extrapolation of $D_{2}(E)$ to the thermodynamic limit.
@@ Appendix A Sierpiński Gasket Geometry
with $N_{g}(1)=9$ active sites in the $g=1$ building block (as illustrated in the main text, MT Fig. 1).
@@ Appendix B On the Hierarchy of Spectral Gaps
the (shifted and scaled) energy levels $x_{g}$ of generation $g$ are related to those of generation g1g-1 by the recurrence relation [82]:
@@
with $\gamma=15/4$ for the standard SG case. This iterative process naturally defines a hierarchy: an initial energy interval containing the spectrum is recursively divided into subintervals, leading to the formation of a self-similar structure of bands and gaps. The system described by Eq. (16) can be approximated by a piece-wise linear model, which allows the retrieval of the full distribution of all gap lengths, as reported in [66].
Appendix C Additional Results for the Representative Extended, Localized, and Non-Ergodic Multifractal States
C.1 Amplitude Profiles
The spatial distributions of the representative extended, localized, and non-ergodic multifractal (NEM) states, as analyzed in the MT, are visualized through their wavefunction intensity profiles, in Fig. 10.
Figure 10: Intensity profiles of the representative ergodic ($E\simeq-3.667$), localized ($E\simeq-2.184$), and NEM ($E\simeq 1.509$) states. Red represents high intensity, black negligible intensity. For the sake of simplicity, the defects that reshape the triangular lattice into a SG are shown as voids. In the ergodic phase, the intensity is homogeneously distributed across all sites, indicating delocalization. The localized state exhibits a strong concentration of intensity in a small region, in this case near the corner. The NEM state has a heterogeneous intensity profile with strong fluctuations that differs from either extended or localized states, and repeats across all scales.
C.2 Spatial Decay Profile of the Localized State
To quantitatively characterize the spatial confinement of the representative localized state ($E\simeq-2.184$), we fit its wavefunction intensity profile, $|\psi(r)|^{2}$, against exponential and power-law decay models as a function of distance $r$ from its peak. To distinguish between the models, we perform a linear regression on the data using both semi-log and log-log scales. A comparison of the resulting coefficients of determination ($R^{2}$) shows a clear preference for the exponential model. For statistical robustness, the fit presented in Fig. 11 aggregates data from generations 6 to 10.
Figure 11: Decay profile analysis for the representative localized state at $E\simeq-2.184$. The main panel (semi-log scale) and inset (log-log scale) test for exponential and power-law decay, respectively. The resulting coefficients of determination are $R^{2}=0.970$ for the exponential fit and $R^{2}=0.844$ for the power-law fit, favoring the former.
C.3 Supplemental Analyses of the Singularity Spectrum
We present two supplemental analyses of the singularity spectrum. The first, presented in Fig. 12, is a systematic cross-check using the direct histogram method as defined in the MT (Methods, Eq. (9)). The second, in Fig. 13, is a study of the finite-size convergence for the NEM state using a sliding window.
Figure 12: Singularity spectrum $f(\alpha)$ calculated using the direct histogram method for the representative states: Extended (GS, $E\simeq-3.667$, left), Non-Ergodic Multifractal (NEM, $E\simeq 1.509$, middle), and Localized (LOC, $E\simeq-2.184$, right). The spectra for generations $g=7$ to $g=10$ provide visual support for our claims, and the distinct character of each state becomes more definite as the system size grows. Although finite, these data provide a reliable basis for the extrapolation methods used in the main text.
Figure 13: Finite-size scaling of the singularity spectrum $f(\alpha)$ for the NEM state at $E\simeq 1.509$. Each curve represents a sliding window of three consecutive system generations. The inset table lists the corresponding spectrum width ($\Delta\alpha$) and peak position ($\alpha_{\text{peak}}$). Crucially, neither the peak nor the width converges towards the expected ergodic value, a behavior that reinforces the state’s non-ergodic nature.
@@ Appendix D Other examples of NEMs
In this section, we strengthen our findings on the emergence of NEM states, confirming that they are not numerical anomalies. We present additional data for several representative multifractal states at different energies, all analyzed following the methodologies detailed in the MT. Specifically, their thermodynamic mass exponents $\tau(E)$ and the related singularity spectra $f(\alpha)$ are presented in Fig. 14. Furthermore, the spectral density-density correlation functions $\mathcal{F}(\omega,E)$ for these same states are displayed in Fig. 15, and their wavefunction intensity profiles are shown in Fig. 16.
Figure 14: Multifractal analysis of Non-Ergodic Multifractal (NEM) states. Left panel: Thermodynamic mass exponents $\tau(E)$ as a function of the moment order $q$ for three NEM states at energies $E\simeq-1.481$ (red circles), $E\simeq-1.287$ (orange squares), and $E\simeq 0.752$ (blue triangles). The dashed line represents the theoretical expectation for GOE statistics ($D_{q}=1$, implying $\tau(q)=q-1$), while the dotted line indicates behavior typical of localized states ($\tau(q)\simeq 0$ for $q>0$). Right panel: The singularity spectra $f(\alpha)$ for the same NEM states, illustrating their different strength of multifractality.
Figure 15: Frequency correlator $\mathcal{F}(\omega,E)$ for three other representative non-ergodic multifractal (NEM) states on a generation $g=6$ SG, compared to the Gaussian Orthogonal Ensemble (GOE) benchmark (dashed line). The states correspond to energies $E\simeq-1.481$ (red), $E\simeq-1.287$ (blue), and $E\simeq 0.752$ (orange). The distinct power-law fall-off of each correlator distinguishes each state.
Figure 16: Intensity profiles of different types of Non-Ergodic Multifractal (NEM) states. The color scale indicates $|\psi|^{2}$ intensity, from negligible (black regions) to high (red regions).
Top row: Intensity distributions for the NEM states detailed in Sec. D. Energies are (from left to right): $E\simeq-1.481$, $E\simeq-1.287$, and $E\simeq 0.752$. Bottom row: Other wavefunction structures at energies (from left to right): $E\simeq-3.554$, $E\simeq-1.018$, and $E\simeq-0.498$. In these supplementary examples, some states are made with “pieces of multifractals” and display properties of either insulators or metals [55], while others have spatial features reminiscent of bulk and edge modes previously identified in fractal geometries [34, 35].
@@ Appendix E Characterization of Weakly Ergodic and Special Multifractal Eigenstates
The SG eigenspectrum hosts states with features that go beyond the exemplary cases classified in the MT. This section presents a study of examples of weakly ergodic states [83] (specifically at energies $E\simeq 0.000$ and $E\simeq 2.000$) and also examines the state at $E\simeq 1.000$ that shows special properties. This latter state is remarkable for its robustness and potential for quantum control, as discussed in the MT, warranting a more detailed investigation of its characteristics here. For these states, we employ the same finite-size scaling methodology for mass exponents $\tau_{q}(E)$ and additionally examine their histogram singularity spectra, wavefunction amplitudes, and statistical properties (such as skewness).
E.1 Multifractal Analysis
The finite-size scaling of the effective mass exponents $\tau_{q}(E,N)$ – calculated as per Eq. (6) from the MT – is performed for system sizes ranging from generation $g=5$ to $g=10$ and presented in Fig. 17. The thermodynamic values $\tau_{q}(E)$ are displayed in the left panel of Fig. 18.
@@ E.1 Multifractal Analysis
Figure 17: Finite-size scaling extrapolation of mass exponents for representative states of energies: $E\simeq 0.000,E\simeq 1.000,E\simeq 2.000$. The thermodynamic value, $\tau_{q}(E)$, is obtained as the $y$-intercept of a linear fit (dashed lines) to the data points from SG generations $g=5$ through $g=10$.
Here, one can observe that while for small moment orders $q$ the mass exponents $\tau_{q}(E)$ for these states tend to follow the $\tau_{q}\approx q-1$ behavior characteristic of extended systems, deviations emerge at larger $q$. Although the thermodynamic fractal dimension $D_{2}$ of these states is high ($0.85\lesssim D_{2}\lesssim 0.95$), there are deviations—in particular for $E\simeq 1.000$—in $\tau_{q}(E)$ at $q>2$, indicating a departure from simple ergodicity and calling for a more detailed examination. The corresponding singularity spectra are displayed in the right panel of Fig. 18 and in Fig. 19.
Figure 18: Left panel: Mass exponents $\tau_{q}(E)$ as a function of moment order qqfor representative states at energies: $E\simeq 0.000$ (red circles), $E\simeq 1.000$ (orange squares), $E\simeq 2.000$ (blue triangles). Statistical errors are included; where not visible, these error bars are smaller than the marker size. The dashed black line indicates the behavior of ideal extended states ($\tau_{q}=q-1$), while the dotted black line represents ideal localized states ($\tau_{q}=0$).
Right panel: The singularity spectra for these states, obtained via Legendre transform, showing a pronounced parabolic shape.
Figure 19: Singularity spectrum $f(\alpha)$ calculated using the direct histogram method for the states at $E\simeq 0.000$ (left), $E\simeq 1.000$ (middle), and $E\simeq 2.000$ (right). Data are shown for generations $g=7$ to $g=10$. The histograms for the state $E\simeq 2.000$ suggest a weakly ergodic character, with the peak near $\alpha\simeq 1$. The states at $E\simeq 0.000$ and $E\simeq 1.000$ display unconventional, comb-like spectra, pointing to their unique spatial distribution across a few disjoint sets of sites.